4.The media often claims that high-income Canadians pay the same or lower taxes than low-income Canadians. Is this a valid claim? Why or why not? Compare a Canadian earning $40,000 per year with a Canadian earning$80,000 per year. Would the higher-income Canadian pay the same amount of tax? Twice as much tax? Or more than twice as much tax? Explain. Explain why marginal tax rates in Canada are higher than average tax rates.

Use the personal opinion to answer,thanks.
Answer #1

For a Canadian earning $40000 per year,Marginal tax rate would be 15%

For a Canadian earning more than $40000 to $80000,Marginal tax rate would be 20.5%

On $40000- taxable income = 15% of 40000 = 6000

Average tax rate = tax/taxable income = 6000/40000 = 15%

On $80000- taxable income = 15% of 40000+ 20.5% of (80000-40000) = 6000+8200 = 14200

Average tax rate = 14200/80000 = 17.75%

Therefore the Canadian earning more than $40000 is paying more than the person earning $40000 which is more than double.